OpenStudy (anonymous):

Simplify 11 to the negative 4th power over 11 to the 8th power. 1 over 11 to the 12th power 1 over 11 to the 4th power 114 1112

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ \frac{11^{-4} }{11^8} \]

OpenStudy (anonymous):

yup

OpenStudy (phi):

can you rewrite the 11^-4 so we get a +4 ? (remember how to do that ?)

OpenStudy (anonymous):

yes we flip

OpenStudy (anonymous):

i think it would be 1/11^4

OpenStudy (phi):

yes. so we now have \[ \frac{1}{11^4} \cdot \frac{1}{11^8} = \frac{1}{11^4\cdot 11^8}\]

OpenStudy (anonymous):

yup

OpenStudy (phi):

11^4 means 11* 11*11*11 11^8 means 11*11*11*11* 11*11*11*11 what do we get when we multiply that all together?

OpenStudy (anonymous):

121

OpenStudy (phi):

?

OpenStudy (anonymous):

you said when we multiply it together?

OpenStudy (phi):

I don't mean multiply it out and get a number. just show them being multiplied like this 11* 11*11*11 * 11* 11*11*11* 11* 11*11*11

OpenStudy (anonymous):

oh woops lol sorry

OpenStudy (phi):

people use exponents to write that in short-hand. can you ?

OpenStudy (anonymous):

umm idk what short hand is

OpenStudy (phi):

we have 11 times itself how many times ? 11* 11*11*11 * 11* 11*11*11* 11* 11*11*11

OpenStudy (anonymous):

12 times

OpenStudy (phi):

and how do we show that using exponents?

OpenStudy (anonymous):

11^12?

OpenStudy (phi):

yes. notice that is easier to type than 11* 11*11*11 * 11* 11*11*11* 11* 11*11*11

OpenStudy (anonymous):

way easier

OpenStudy (anonymous):

ohhh thats the answer to

OpenStudy (phi):

so here is what you have \[ \frac{11^{-4} }{11^8}= \frac{ 1}{11^{4}\cdot 11^8}= \frac{1}{11^{12}} \]

OpenStudy (anonymous):

yup

OpenStudy (phi):

you should notice that 11^4 * 11^8 means 11* 11*11*11 * 11* 11*11*11* 11* 11*11*11 but it is faster to say: 4 11's and 8 11's is 12 11's in other words, notice the short cut: 4+8 =12 11^4 * 11^8 = 11^12

OpenStudy (anonymous):

mmmhhh

OpenStudy (anonymous):

it takes up space ifyou write it out also

OpenStudy (phi):

the thing you should try to remember is when multiplying, you can add the exponents. that is the short cut they want you to learn.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the answer is 1/11^12

OpenStudy (phi):

yes and if they were being trick the answer could also be 11^(-12)

OpenStudy (anonymous):

yuppp

OpenStudy (anonymous):

thanks again and i have like 2 or 3 more if you can still help it would be great

OpenStudy (phi):

new post

OpenStudy (anonymous):

ok